Overcoming a Head Start Review Part A Cars A and B are racing each other along the same straight road in the following manner Car A has a head start and is a distance DA beyond the starting line at t-0. The starting line is at z-0. Car A travels at a constant speed vA. Car B starts at the starting line but has a better engine than Car A, and thus Car B travels at a constant speed vp, which is greater than vA How long after Car B started the race will Car B catch up with Car A? Express the time in terms of given quantities View Available Hint(s) tcatch - Submit Part B How far from Car Bs starting line will the cars be when Car B passes Car A? Express your answer in terms of known quantities. (You may use teatch as well.) View Available Hint(s) Submit Provide Feedback
Answer #1

A)

let car B catch car A after time "tcatch"

VB = speed of car B

VA = speed of car A

XAi = initial position of car A = DA

XA = position of car A after time "tcatch" = ?

Position of car A after time "tcatch" is given as

XA = XAi + VA tcatch

XA = DA + VA tcatch                                                    eq-1

XBi = initial Position of car B = 0

XB = Position of car B after time "tcatch" = ?

Position of car B after time "tcatch" is given as

XB = XBi + VB tcatch

XB = 0+ VB tcatch                                                    eq-2

for the car A to catch car B ,

XA = XB

DA + VA tcatch = VB tcatch

VB tcatch - VA tcatch = DA

tcatch = DA / (VB - VA )

b)

position of car B is given as

XB = 0+ VB tcatch

XB = VB (DA / (VB - VA ))

XB = VB DA / (VB - VA )

dpass = VB DA / (VB - VA )